# The coarse Baum-Connes conjecture with filtered coefficients and product metric spaces

### v1

@@ Abstract.
Inspired by the quantitative $K$-theory, in this paper, we introduce the coarse Baum-Connes conjecture with filtered coefficients which generalizes the original conjecture. The are two advantages for the conjecture with filtered coefficients. Firstly, the routes toward the coarse Baum-Connes conjecture also work for the conjecture with filtered coefficients. Secondly, the class of metric spaces that satisfy the conjecture with filtered coefficients is closed under products and yet it is not known for the original conjecture. As an application, we show some new examples of product metric spaces for the coarse Baum-Connes conjecture.
The author was supported by NSFC12171156, 12271165, 12301154.
@@ 1. Introduction
The coarse Baum-Connes conjecture (cf. [8][19]) is dated back to the celebrated Atiyah-Singer index theorem which gives a formula by using topological index to compute the Fredholm index of an elliptic differential operator on a closed manifold. For a Riemannian manifold $M$ and an elliptic differential operator $D$ on $M$, Roe defined a higher index of $D$ which lives in the $K$-theory of a $C^{\ast}$-algebra, now called Roe algebra of $M$ (cf. [16]). Let us recall its definition. Choose a locally finite net $N_{M}$ in $M$, the Roe algebra of $M$ is defined to be the norm closure of the $\ast$-algebra consisting of all $(N_{M}\times N_{M})$-matrices $T$ with coefficients in $\mathfrak{K}(H)$ satisfying that $T_{x,y}=0$ when $d(x,y)>R$ for some $R>0$, where $\mathfrak{K}(H)$ is the algebra of compact operators on a Hilbert space $H$. The coarse Baum-Connes conjecture assert that the assembly map from the coarse $K$-homology of $M$ to the $K$-theory of Roe algebra of $M$ is an isomorphism (see Conjecture 3.23). Thus, the coarse Baum-Connes conjecture provides a topological algorithm for higher indexes. Moreover, the conjecture has important applications to the Novikov conjecture and the existence of Riemannian metric on a manifold with positive scalar curvature (cf. [9][18]).
There are three routes toward the coarse Baum-Connes conjecture. Roughly speaking, the first one is algebraic, in [21], Yu invented the quantitative $K$-theory to verify the conjecture holds for metric spaces with finite asymptotic dimension. The second one is geometric, in [22], Yu used the Dirac-dual-Dirac method to prove that the conjecture holds for metric spaces which admit a coarse embedding into Hilbert space. And the third one is topological, in [8], Higson and Roe using coarse homotopy and open cones to show the conjecture is true for non-positively curved metric spaces. On the other hand, there exist counterexamples to the coarse Baum-Connes conjecture (cf. [7][12]).
In [15], Oyono-Oyono and Yu systemized the quantitative $K$-theory for filtered $C^{\ast}$-algebras as defined below.
@@
([15, Definition 1.1]) Let $A$ be a $C^{\ast}$-algebra. A filtration of $A$ is a family of closed linear subspaces $(A_{r})_{r>0}$ such that
@@
If $A$ is unital, we also require the unit of $A$ belongs to $A_{r}$ for any $r>0$. A $C^{\ast}$-algebra equipped with a filtration is called a filtered $C^{\ast}$-algebra.
As a refinement of the $K$-theory, the quantitative $K$-theory with order $r$ of $A$ considers homotopy equivalent classes of almost-projections or almost-unitaries in $A_{r}$.
Inspired by the quantitative $K$-theory, we define the Roe algebra of $M$ with filtered coefficients in $A$ to be the norm closure of the $\ast$-algebra consisting of all $(N_{M}\times N_{M})$-matrices $T$ with coefficients in $\mathfrak{K}(H)\otimes A_{r}$ satisfying that $T_{x,y}=0$ when $d(x,y)>R$ for some $r,R>0$ (see Definition 3.2). And we introduce the coarse Baum-Connes conjecture with filtered coefficients to compute the $K$-theory of Roe algebras with filtered coefficients by using of the $K$-theory of localization algebras with filtered coefficients (see Conjecture 3.23). When $A=\mathbb{C}$ with the trivial filtration $A_{r}=\mathbb{C}$, the coarse Baum-Connes conjecture with filtered coefficients in $\mathbb{C}$ is exactly the original conjecture. Moreover, the above-mentioned three ways to the original conjecture also work for the coarse Baum-Connes conjecture with filtered coefficients (see Section 5).
In order to explore the coarse Baum-Connes conjecture with filtered coefficients for product metric spaces, we introduce a notion of localization algebras along one direction with filtered coefficients as a bridge (see Definition 4.3). Then we obtain the following result (see also Theorem 4.19).
@@
    the coarse Baum-Connes conjecture with filtered coefficients in $C^{\ast}_{f}(Y,A)$ holds for $X$;
@@
    the coarse Baum-Connes conjecture with filtered coefficients in $C^{\ast}_{f}(\square,A)$ holds for $Y$, where $\square$ is the separated coarse union of a sequence of single points.
In particular, if the coarse Baum-Connes conjecture with filtered coefficients holds for $X$ and $Y$, then the coarse Baum-Connes conjecture with filtered coefficients holds for $X\times Y$.
The above theorem is a coarse analogue of H. Oyono-Oyono’s result in [14] stating that the Baum-Connes conjecture with coefficients is closed under the direct product of groups.
In [5], Fukaya and Oguni proved that the coarse Baum-Connes conjecture holds for product metric spaces of certain groups by studying the coronae of products. In generally, if $X$ and $Y$ satisfy the coarse Baum-Connes conjecture, it is not known whether the conjecture is true or not for $X\times Y$. Take $A=\mathbb{C}$ in Theorem 1.2, we have the following corollary (see also Corollary 4.20) which not only recovers Fukaya and Oguni’s result but also provides new examples of products satisfying the conjecture (see Section 5).
@@
    $X$ satisfies the coarse Baum-Connes conjecture with filtered coefficients in $C^{\ast}(Y)$;
@@
    $Y$ satisfies the coarse Baum-Connes conjecture with coefficients in $\ell^{\infty}(\mathbb{N},\mathfrak{K}(H))$.
Then the coarse Baum-Connes conjecture holds for $X\times Y$.
@@
Besides, the coarse Baum-Connes conjecture with filtered coefficients is also applied to the quantitative coarse Baum-Connes conjecture which is a refinement of the original conjecture, we will explore this connection in the subsequent paper (cf. [23]).
The paper is organized as follows. In Section 2, we recall some ingredients from the $K$-theory of $C^{\ast}$-algebras. In Section 3, we introduce Roe algebras with filtered coefficients, localization algebras with filtered coefficients and discuss their properties, finally, we formulate the coarse Baum-Connes conjecture with filtered coefficients. In Section 4, we give the coarse Baum-Connes conjecture along one direction for product metric spaces, then build a connection between it and the conjecture with filtered coefficients, finally, we prove Theorem 1.2. In Section 5, we apply our results to a number of metric spaces .
@@ 2. Preliminaries
In this section, we recall three important techniques for computing the $K$-theory of $C^{\ast}$-algebras. The main reference is Willett and Yu’s book [18].
@@
[18, Proposition 2.7.5] Let $\alpha:C\rightarrow D$ be a $\ast$-homomorphism and $v$ be a partial isometry in the multiplier algebra of $D$ satisfying that $\alpha(c)v^{\ast}v=\alpha(c)$ for all $c\in C$. Define $\beta:C\rightarrow D$ by $\beta(c)=v\alpha(c)v^{\ast}$. Then $\alpha$ and $\beta$ induce the same morphism on $K$-theory.
@@
[18, Definition 2.7.11] A $C^{\ast}$-algebra $B$ is called to be quasi-stable if for all positive integer $n$ there exists an isometry $v$ in the multiplier algebra of $M_{n}(A)$ such that $vv^{\ast}$ is the matrix unit $e_{11}$.
@@
[18, Lemma 12.4.3] Let $B$ be a quasi-stable $C^{\ast}$-algebra and let $C_{ub}([0,\infty),B)$ be the $C^{\ast}$-algebra consisting of all bounded, uniformly continuous functions from $[0,\infty)$ to $B$, then the evaluation at zero map
@@
The following Mayer-Vietoris six-term exact sequence is a key technique to compute the $K$-theory of $C^{\ast}$-algebras by their ideals, for the proof, please refer to [18, Proposition 2.7.15] or [10].
@@
Now, Let us fix some notations in the rest of this paper. Let $(X,d)$ be a proper metric space (the properness means every closed ball in $X$ is compact). Since that proper metric spaces are separable, thus we can choose a countable dense subset $Z_{X}$ in $X$. And let $A$ be a filtered $C^{\ast}$-algebra acting on a Hilbert space $H_{A}$. Fixed $H$ as a separable Hilbert space.
@@ 3.1. Roe algebras with filtered coefficients
The Roe algebra was introduced by Roe in [17] motivated by the index theory on open manifolds [16]. In this subsection, we introduce a notion of Roe algebra with filtered coefficients in a filtered $C^{\ast}$-algebra.
@@
Let $X$, $Y$ be two proper metric spaces and $Z_{X}$, $Z_{Y}$ be two countable dense subsets of $X$, $Y$, respectively. Let $\chi$ be the characteristic function.
@@
    For a bounded operator $T:\ell^{2}(Z_{X})\otimes H\otimes H_{A}\rightarrow\ell^{2}(Z_{Y})\otimes H\otimes H_{A}$, the support of $T$, denoted by $\operatorname{supp}(T)$, is defined to be the set consisting of all $(x,y)\in X\times Y$ satisfying that $(\chi_{V}\otimes I\otimes I)T(\chi_{U}\otimes I\otimes I)\neq 0$ for any open neighborhoods $U$ and $V$ of $x$ and $y$, respectively.
@@
    For a bounded operator $T$ on $\ell^{2}(Z_{X})\otimes H\otimes H_{A}$, the propagation of $T$, denoted by $\operatorname{prop}(T)$, is defined to be $\sup\{d(x,y):(x,y)\in\operatorname{supp}(T)\}$.
@@
  • there exists $s>0$ such that for any compact subset $K\subseteq X$, the operators $(\chi_{K}\otimes I\otimes I)T$ and $T(\chi_{K}\otimes I\otimes I)$ are in $\mathfrak{K}(\ell^{2}(Z_{X})\otimes H)\otimes A_{s}$, where $\mathfrak{K}(\ell^{2}(Z_{X})\otimes H)$ is the algebra of compact operators on $\ell^{2}(Z_{X})\otimes H$.
@@
If $A$ is a $C^{\ast}$-algebra is equipped with the trivial filtration, namely, $A_{r}=A$ for any $r>0$, then $C_{f}^{\ast}(X,A)$ is the Roe algebra of $X$ with coefficients in $A$, also denoted by $C^{\ast}(X,A)$. In particular, when $A=\mathbb{C}$, the $C^{\ast}$-algebra $C_{f}^{\ast}(X,\mathbb{C})$ or $C^{\ast}(X)$ for short is so called the Roe algebra of $X$ (cf. [17]).
@@
Two proper metric spaces $X$ and $Y$ are called to be coarsely equivalent, if there exists two coarse maps $g:X\rightarrow Y$, $h:Y\rightarrow X$ and a constant $c>0$ such that
$\max\{d(hg(x),x),d(gh(y),y)\}\leq c,$
@@
For a coarse map $g:X\rightarrow Y$ and any $\delta>0$. Choose a disjoint Borel cover $(U_{i})_{i\in\mathbb{N}}$ of $Y$ with diameter less than $\delta$ and every $U_{i}$ has non-empty interior, then there exists a family of isometries $V_{i}:\ell^{2}(g^{-1}(U_{i})\cap Z_{X})\otimes H\rightarrow\ell^{2}(U_{i}\cap Z_{Y})\otimes H$. Let
$V_{g}=\bigoplus_{i\in\mathbb{N}}V_{i}\otimes I:\ell^{2}(Z_{X})\otimes H\otimes H_{A}\rightarrow\ell^{2}(Z_{Y})\otimes H\otimes H_{A}.$
@@
Then we have $K^{\prime}\subseteq g^{-1}(\overline{B_{\delta}(K)})$ since $d(g(x),y)<\delta$ for any $(y,x)\in\operatorname{supp}(V^{\ast}_{g})$, where $\overline{B_{\delta}(K)}=\{y\in Y:d(y,K)\leq\delta\}$ which is compact by the properness of $Y$. Thus $K^{\prime}$ is a compact subset of $X$ by the properness of $g$. Moreover, we have that $V_{g}TV^{\ast}_{g}(\chi_{K}\otimes I\otimes I)=V_{g}T(\chi_{K^{\prime}}\otimes I\otimes I)V^{\ast}_{g}(\chi_{K}\otimes I\otimes I)$ which implies $V_{g}TV^{\ast}_{g}(\chi_{K}\otimes I\otimes I)\in\mathfrak{K}(\ell^{2}(Z_{X})\otimes H)\otimes A_{s}$ since $T(\chi_{K^{\prime}}\otimes I\otimes I)\in\mathfrak{K}(\ell^{2}(Z_{X})\otimes H)\otimes A_{s}$. Similarly, we can obtain the similar result for $(\chi_{K}\otimes I\otimes I)V_{g}TV^{\ast}_{g}$. In conclusion, $ad_{g}$ is well-defined.
Finally, if there exists another isometry $V^{\prime}_{g}:\ell^{2}(Z_{X})\otimes H\otimes H_{A}\rightarrow\ell^{2}(Z_{Y})\otimes H\otimes H_{A}$ satisfying the condition (3.1). Then by similar arguments as above, $V_{g}V^{\ast}_{g}$, $V_{g}^{\prime}V^{\prime\ast}_{g}$, $V^{\prime}_{g}V^{\ast}_{g}$ are all multipliers of $C_{f}^{\ast}(Y,A)$. And
$\begin{pmatrix}0&0\\ 0&V^{\prime}_{g}TV^{\prime\ast}_{g}\end{pmatrix}=\begin{pmatrix}I-V_{g}V^{\ast}_{g}&V_{g}V^{\prime\ast}_{g}\\ V^{\prime}_{g}V^{\ast}_{g}&I-V^{\prime}_{g}V^{\prime\ast}_{g}\end{pmatrix}\begin{pmatrix}V_{g}TV^{\ast}_{g}&0\\ 0&0\end{pmatrix}\begin{pmatrix}I-V_{g}V^{\ast}_{g}&V_{g}V^{\prime\ast}_{g}\\ V^{\prime}_{g}V^{\ast}_{g}&I-V^{\prime}_{g}V^{\prime\ast}_{g}\end{pmatrix}$
which implies that $ad^{\prime}_{g}$ and $ad_{g}$ induce the same homomorphism on $K$-theory by Lemma 2.1 because their images are unitarily equivalent by a unitary multiplier. ∎
If $g$ is the identity map on $X$, then we can choose $V_{g}$ to be a unitary, thus we obtain the following corollary.
@@
$V_{g}=\bigoplus_{i\in\mathbb{N}}V_{i}\otimes I:\ell^{2}(Z_{X})\otimes H\otimes H_{A}\rightarrow\ell^{2}(Z_{Y})\otimes H\otimes H_{A}.$
Then $V_{g}$ is a unitary and we have that
@@
The algebra $C_{f}^{\ast}(X,A)$ is a filtered $C^{\ast}$-algebra equipped with a filtration $C_{f}^{\ast}(X,A)_{r}$ consisting of all operators $T\in C_{f}^{\ast}(X,A)$ such that $\operatorname{prop}(T)\leq r$ and $(\chi_{K}\otimes I\otimes I)T,T(\chi_{K}\otimes I\otimes I)\in\mathfrak{K}(\ell^{2}(Z_{X})\otimes H)\otimes A_{r}$ for any compact subset $K\subseteq X$.
@@
$\{(x,y)\in X\times X:\>\>\textrm{there exists $z\in X$ such that $(x,z)\in\operatorname{supp}(S)$, $(z,y)\in\operatorname{supp}(T)$}\}$
@@
Besides, since that $(A_{r})_{r>0}$ is a filtration of $A$, thus $(C_{f}^{\ast}(X,A)_{r})_{r>0}$ is a filtration of $C_{f}^{\ast}(X,A)$. ∎

### v2

@@ Abstract.
Inspired by the quantitative $K$-theory, in this paper, we introduce the coarse Baum-Connes conjecture with filtered coefficients which generalizes the original conjecture. There are two advantages for the conjecture with filtered coefficients. Firstly, the routes toward the coarse Baum-Connes conjecture also work for the conjecture with filtered coefficients. Secondly, the class of metric spaces that satisfy the conjecture with filtered coefficients is closed under products and yet it is unknown for the original conjecture. As an application, we discover some new examples of product metric spaces for the coarse Baum-Connes conjecture.
This work was supported by NSFC (Nos. 12171156, 12271165, 12301154) and the Fundamental Research Funds for the Central Universities (No. 1301032574).
@@ 1. Introduction
The coarse Baum-Connes conjecture (cf. [9][22]) can be dated back to the celebrated Atiyah-Singer index theorem which gives a formula by using the topological index to compute the Fredholm index of an elliptic differential operator on a closed manifold. For an elliptic differential operator on an open Riemannian manifold, Roe defined a higher index which lives in the $K$-theory of a $C^{\ast}$-algebra, now called the Roe algebra (cf. [18]). Let us recall its definition for a locally finite metric space $X$. Let $\mathfrak{K}(H)$ be the algebra of compact operators on a Hilbert space $H$. The Roe algebra of $X$, denoted by $C^{\ast}(X)$, is defined to be the norm closure of the $\ast$-algebra consisting of all $(X\times X)$-matrices $T$ with coefficients in $\mathfrak{K}(H)$ satisfying $T_{x,y}=0$ when $d(x,y)>R$ for some $R>0$.
The coarse Baum-Connes conjecture for $X$ asserts that the assembly map from the coarse $K$-homology of $X$ to the $K$-theory of the Roe algebra of $X$ is an isomorphism (see Conjecture 3.23). The left-hand side of the conjecture is topological and computable. Thus, the coarse Baum-Connes conjecture provides a topological algorithm for higher indexes of elliptic differential operators on Riemannian manifolds. Moreover, the conjecture has important applications to the Novikov conjecture on the homotopy invariance of higher signatures for closed manifolds, the Gromov-Lawson-Rosenberg conjecture concerning the existence of Riemannian metrics with positive scalar curvature on a manifold and Gromov’s zero-in-the-spectrum conjecture about the spectrum of the Laplacian operator on a Riemannian manifold (cf. [10][19][21]).
There are three routes toward the coarse Baum-Connes conjecture. Roughly speaking, the first one is algebraic. In [24], Yu introduced the quantitative $K$-theory to verify that the conjecture holds for metric spaces with finite asymptotic dimension. The second one is geometric. In [25], Yu applied the Dirac-dual-Dirac method to prove that the conjecture holds for metric spaces which admit a coarse embedding into Hilbert space. The third one is topological. In [9], Higson and Roe employed the coarse homotopy and open cones to show that the conjecture is true for non-positively curved metric spaces. On the other hand, there exist some counterexamples to the coarse Baum-Connes conjecture, such as large spheres and certain expanders or asymptotic expanders (cf. [8][13][20][24]).
In [16], Oyono-Oyono and Yu generalized the quantitative $K$-theory to all filtered $C^{\ast}$-algebras as defined below.
@@
([16, Definition 1.1]) Let $A$ be a $C^{\ast}$-algebra. A filtration of $A$ is a family of closed linear subspaces $(A_{r})_{r>0}$ such that
@@
If $A$ is unital, we also require the unit of $A$ belongs to $A_{r}$ for each $r>0$. A $C^{\ast}$-algebra equipped with a filtration is called a filtered $C^{\ast}$-algebra.
As a refinement of the $K$-theory, the quantitative $K$-theory with order $r$ of $A$ is defined by homotopy equivalent classes of almost-projections (or almost-unitaries) in $n\times n$-matrices with coefficients in $A_{r}$ for all $n\in\mathbb{N}$.
Inspired by the quantitative $K$-theory, in this paper, we introduce the notion of the Roe algebra with filtered coefficients as below (see Definition 3.2 for the general case).
Definition 1.2.
Let $X$ be a locally finite metric space and $A$ be a filtered $C^{\ast}$-algebra. The Roe algebra of $X$ with filtered coefficients in $A$, denoted by $C^{\ast}_{f}(X,A)$, is defined to be the norm closure of the $\ast$-algebra consisting of all $(X\times X)$-matrices $T$ with coefficients in $\mathfrak{K}(H)\otimes A_{r}$ satisfying $T_{x,y}=0$ when $d(x,y)>R$ for some $r,R>0$.
Then we introduce the coarse Baum-Connes conjecture with filtered coefficients to compute the $K$-theory of Roe algebras with filtered coefficients by using the $K$-theory of localization algebras with filtered coefficients (see Conjecture 3.23). When $A=\mathbb{C}$, the coarse Baum-Connes conjecture with filtered coefficients in $\mathbb{C}$ is exactly the original conjecture. Moreover, the above-mentioned three ways to the original conjecture also work for the coarse Baum-Connes conjecture with filtered coefficients (see Section 5).
In order to connect the coarse Baum-Connes conjecture with filtered coefficients for product metric spaces and the conjecture with filtered coefficients for their factors, we introduce a notion of localization algebras along one direction with filtered coefficients as a bridge (see Definition 4.3). Then by a Mayer-Vietoris argument, we obtain the following main result (see also Theorem 4.21).
@@
    $X$ satisfies the coarse Baum-Connes conjecture with filtered coefficients in the Roe algebra $C^{\ast}_{f}(Y,A)$;
@@
    $Y$ satisfies the coarse Baum-Connes conjecture with filtered coefficients in the uniform product $\prod^{u}_{\mathbb{N}}(A\otimes\mathfrak{K}(H))$.
Then $X\times Y$ satisfies the coarse Baum-Connes conjecture with filtered coefficients in $A$.
In particular, if $X$ and $Y$ satisfy the coarse Baum-Connes conjecture with filtered coefficients, then $X\times Y$ satisfies the coarse Baum-Connes conjecture with filtered coefficients.
Theorem 1.3 is a coarse analogue of Oyono-Oyono’s result in [15, Corollary 7.12] stating that the Baum-Connes conjecture with coefficients is closed under the direct product of groups.
In [6], Fukaya and Oguni proved that the coarse Baum-Connes conjecture holds for product metric spaces of certain geometric groups by studying the coronae of products. In generally, if $X$ and $Y$ satisfy the coarse Baum-Connes conjecture, it is unknown whether the conjecture is true or not for $X\times Y$. Taking $A=\mathbb{C}$ in Theorem 1.3, we have the following corollary (see also Corollary 4.23), which not only recovers Fukaya and Oguni’s result but also provides some new examples of products satisfying the conjecture (see Section 5).
@@
    $X$ satisfies the coarse Baum-Connes conjecture with filtered coefficients in the Roe algebra $C^{\ast}(Y)$;
@@
    $Y$ satisfies the coarse Baum-Connes conjecture with coefficients in the direct product $\prod_{\mathbb{N}}\mathfrak{K}(H)$.
Then $X\times Y$ satisfies the coarse Baum-Connes conjecture.
@@
Corollary 1.4 is a general result for the coarse Baum-Connes conjecture of product metric spaces since we do not require any geometric conditions on each metric space.
Besides, the coarse Baum-Connes conjecture with filtered coefficients is also applied to the quantitative coarse Baum-Connes conjecture, which is a refinement of the original conjecture. We will explore this connection in a subsequent paper (cf. [26]).
The paper is organized as follows. In Section 2, we recall some ingredients from the $K$-theory of $C^{\ast}$-algebras. In Section 3, we introduce Roe algebras with filtered coefficients, localization algebras with filtered coefficients and the (consistent) coarse Baum-Connes conjecture with filtered coefficients. In Section 4, we introduce the localization algebras along one direction for products to build a connection between the coarse Baum-Connes conjecture with filtered coefficients for products and the conjecture with filtered coefficients for their factors. And we prove Theorem 1.3 in the end of this section. In Section 5, we apply our results to a number of product metric spaces.
@@ 2. Preliminaries
In this section, we shall recall three important techniques for computing the $K$-theory of $C^{\ast}$-algebras. The main reference is Willett and Yu’s book [21].
@@
[21, Proposition 2.7.5] Let $\alpha:C\rightarrow D$ be a $\ast$-homomorphism and $v$ be a partial isometry in the multiplier algebra of $D$ satisfying $\alpha(c)v^{\ast}v=\alpha(c)$ for all $c\in C$. Define $\beta:C\rightarrow D$ by $\beta(c)=v\alpha(c)v^{\ast}$. Then $\alpha$ and $\beta$ induce the same morphism on $K$-theory.
@@
[21, Definition 2.7.11] A $C^{\ast}$-algebra $B$ is called to be quasi-stable if for any positive integer $n$ there exists an isometry $v$ in the multiplier algebra of $M_{n}(B)$ such that $vv^{\ast}$ is the matrix unit $e_{11}$.
@@
[21, Lemma 12.4.3] Let $B$ be a quasi-stable $C^{\ast}$-algebra and $C_{ub}([0,\infty),B)$ be the $C^{\ast}$-algebra consisting of all bounded, uniformly continuous functions from $[0,\infty)$ to $B$. Then the evaluation at zero map
@@
The following Mayer-Vietoris six-term exact sequence is a key technique to compute the $K$-theory of $C^{\ast}$-algebras by their ideals. Please refer to [21, Proposition 2.7.15] or [11] for the proof.
@@
Now, Let us fix some notations in the rest of this paper. Let $(X,d)$ be a proper metric space, namely, every closed ball in $X$ is compact. Since proper metric spaces are separable, thus we can choose a countable dense subset $Z_{X}$ in $X$. Let $A$ be a filtered $C^{\ast}$-algebra acting on a Hilbert space $H_{A}$. Fixed $H$ as a separable Hilbert space. And let $\mathfrak{K}(H^{\prime})$ be the algebra of compact operators on a Hilbert space $H^{\prime}$.
@@ 3.1. Roe algebras with filtered coefficients
The Roe algebra was introduced by Roe in [19] motivated by the index theory on open manifolds [18]. In this subsection, we introduce a notion of the Roe algebra with filtered coefficients in a filtered $C^{\ast}$-algebra.
@@
Let $X$ and $Y$ be two proper metric spaces. Let $Z_{X}$ and $Z_{Y}$ be two countable dense subsets of $X$ and $Y$, respectively. Let $\chi$ be the characteristic function.
@@
    For a bounded operator $T:\ell^{2}(Z_{X})\otimes H\otimes H_{A}\rightarrow\ell^{2}(Z_{Y})\otimes H\otimes H_{A}$. The support of $T$, denoted by $\operatorname{supp}(T)$, is defined to be the set consisting of all $(x,y)\in X\times Y$ satisfying $(\chi_{V}\otimes I\otimes I)T(\chi_{U}\otimes I\otimes I)\neq 0$ for any open neighborhoods $U$ and $V$ of $x$ and $y$, respectively.
@@
    For a bounded operator $T$ on $\ell^{2}(Z_{X})\otimes H\otimes H_{A}$. The propagation of $T$, denoted by $\operatorname{prop}(T)$, is defined to be $\sup\{d(x,y):(x,y)\in\operatorname{supp}(T)\}$.
@@
  • there exists $s>0$ such that $(\chi_{K}\otimes I\otimes I)T$ and $T(\chi_{K}\otimes I\otimes I)$ belong to $\mathfrak{K}(\ell^{2}(Z_{X})\otimes H)\otimes A_{s}$ for any compact subset $K\subseteq X$.
@@
If a $C^{\ast}$-algebra $A$ is endowed with the trivial filtration, namely, $A_{r}=A$ for each $r>0$. Then $C_{f}^{\ast}(X,A)$ is the Roe algebra of $X$ with coefficients in $A$, also denoted by $C^{\ast}(X,A)$. In particular, the $C^{\ast}$-algebra $C_{f}^{\ast}(X,\mathbb{C})$ or $C^{\ast}(X)$ for short, is so-called the Roe algebra of $X$.
@@
Two proper metric spaces $X$ and $Y$ are called to be coarsely equivalent, if there exist two coarse maps $g:X\rightarrow Y$, $h:Y\rightarrow X$ and a constant $c>0$ such that
$\max\{d(hg(x),x),d(gh(y),y)\}\leq c$
@@
For a coarse map $g:X\rightarrow Y$ and $\delta>0$. Choosing a disjoint Borel cover $(U_{i})_{i\in\mathbb{N}}$ of $Y$ with diameter less than $\delta$ and every $U_{i}$ has non-empty interior, then there exists a family of isometries $V_{i}:\ell^{2}(g^{-1}(U_{i})\cap Z_{X})\otimes H\rightarrow\ell^{2}(U_{i}\cap Z_{Y})\otimes H$. Let
$V_{g}=\bigoplus_{i\in\mathbb{N}}V_{i}\otimes I_{H_{A}}:\ell^{2}(Z_{X})\otimes H\otimes H_{A}\rightarrow\ell^{2}(Z_{Y})\otimes H\otimes H_{A}.$
@@
Then we have $K^{\prime}\subseteq g^{-1}(\overline{B_{\delta}(K)})$ since $d(g(x),y)<\delta$ for any $(y,x)\in\operatorname{supp}(V^{\ast}_{g})$, where $\overline{B_{\delta}(K)}=\{y\in Y:d(y,K)\leq\delta\}$ is compact by the properness of $Y$. Thus $K^{\prime}$ is a compact subset of $X$ by the properness of $g$. Moreover, since $T(\chi_{K^{\prime}}\otimes I\otimes I)\in\mathfrak{K}(\ell^{2}(Z_{X})\otimes H)\otimes A_{s}$ and $V_{g}TV^{\ast}_{g}(\chi_{K}\otimes I\otimes I)=V_{g}T(\chi_{K^{\prime}}\otimes I\otimes I)V^{\ast}_{g}(\chi_{K}\otimes I\otimes I)$. Thus, we have $V_{g}TV^{\ast}_{g}(\chi_{K}\otimes I\otimes I)\in\mathfrak{K}(\ell^{2}(Z_{X})\otimes H)\otimes A_{s}$. We can obtain the similar result for $(\chi_{K}\otimes I\otimes I)V_{g}TV^{\ast}_{g}$. In conclusion, $ad_{g}$ is well-defined.
Finally, if there exists another isometry $V^{\prime}_{g}:\ell^{2}(Z_{X})\otimes H\otimes H_{A}\rightarrow\ell^{2}(Z_{Y})\otimes H\otimes H_{A}$ satisfying the condition (3.1). Then by similar arguments as above, $V_{g}V^{\ast}_{g}$, $V_{g}^{\prime}V^{\prime\ast}_{g}$, $V^{\prime}_{g}V^{\ast}_{g}$ are all multipliers of $C_{f}^{\ast}(Y,A)$. And we have
$\begin{pmatrix}0&0\\ 0&V^{\prime}_{g}TV^{\prime\ast}_{g}\end{pmatrix}=\begin{pmatrix}I-V_{g}V^{\ast}_{g}&V_{g}V^{\prime\ast}_{g}\\ V^{\prime}_{g}V^{\ast}_{g}&I-V^{\prime}_{g}V^{\prime\ast}_{g}\end{pmatrix}\begin{pmatrix}V_{g}TV^{\ast}_{g}&0\\ 0&0\end{pmatrix}\begin{pmatrix}I-V_{g}V^{\ast}_{g}&V_{g}V^{\prime\ast}_{g}\\ V^{\prime}_{g}V^{\ast}_{g}&I-V^{\prime}_{g}V^{\prime\ast}_{g}\end{pmatrix},$
which implies that $ad^{\prime}_{g}$ and $ad_{g}$ induce the same homomorphism on $K$-theory by Lemma 2.1. ∎
If $g$ is the identity map on $X$, then we can choose $V_{g}$ to be a unitary. Thus we obtain the following corollary.
@@
$V_{g}=\bigoplus_{i\in\mathbb{N}}V_{i}\otimes I_{H_{A}}:\ell^{2}(Z_{X})\otimes H\otimes H_{A}\rightarrow\ell^{2}(Z_{Y})\otimes H\otimes H_{A}.$
Then $V_{g}$ is a unitary and we have
@@
The Roe algebra $C_{f}^{\ast}(X,A)$ is a filtered $C^{\ast}$-algebra equipped with a filtration $C_{f}^{\ast}(X,A)_{r}$ consisting of all operators $T\in C_{f}^{\ast}(X,A)$ such that $\operatorname{prop}(T)\leq r$ and $(\chi_{K}\otimes I\otimes I)T,T(\chi_{K}\otimes I\otimes I)\in\mathfrak{K}(\ell^{2}(Z_{X})\otimes H)\otimes A_{r}$ for any compact subset $K\subseteq X$.
@@
$\{(x,y)\in X\times X:\>\>\textrm{there exists $z\in X$ such that $(x,z)\in\operatorname{supp}(S)$, $(z,y)\in\operatorname{supp}(T)$}\},$
@@
Besides, since $(A_{r})_{r>0}$ is a filtration of $A$, thus $(C_{f}^{\ast}(X,A)_{r})_{r>0}$ is a filtration of $C_{f}^{\ast}(X,A)$. ∎